Opened 10 years ago

Closed 3 years ago

#323 closed flight processing (fixed)

GB07/12, flight day 082a/2010, Debden

Reported by: benj Owned by: knpa
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2010/flight_data/uk/GB07_12-2010_082a_Debden

Data arrived from ARSF via network transfer on 25th March 2010.

Scientific objective: Monitoring of peat fires and quantification of pyrogenic carbon conversion

Priority: a4l

PI: Martin Wooster

Note original PI on application was Andrew Wilson, but since he's retired effective PI is Martin Wooster.

Sensors:

  • Eagle (07/09/2010) (reprocessed 23/04/2012)
  • Hawk (20/12/2010) (reprocessed 23/04/2012)
  • Leica LIDAR (23/07/2010)
  • RCD (23/07/2010)

Change History (32)

comment:1 Changed 10 years ago by benj

  • Description modified (diff)
  • Summary changed from GB07/12, flight day 082a/2010, Debden burn to GB07/12, flight day 082a/2010, Debden

comment:2 Changed 10 years ago by benj

  • Milestone set to 2010 data processing completion

comment:3 Changed 9 years ago by mark1

removed rcd logs that related to flight 082b Debden

comment:4 Changed 9 years ago by mark1

  • Owner set to mark1
  • Status changed from new to accepted

comment:5 Changed 9 years ago by mark1

Have renamed raw eagle flight line files 4-10 from 082-10-*.* to VNIR082-10-*.*

Also renamed 082-10-1.nav to VNIR082-10-2.nav (there is already a file named VNIR082-10-1.nav)

comment:6 Changed 9 years ago by mark1

Have renamed hawk flight line files 4-10 from 082-10-*.* to SWIR082-10-*.*

comment:7 Changed 9 years ago by mark1

Original RCD images 1-8 are test images and not processed. Images 10 and 11 failed and have been removed from delivery.

comment:8 Changed 9 years ago by mark1

Photograph delivery check complete.

comment:9 Changed 9 years ago by mark1

Typed logsheet has been received from Ops and added to admin dir.

comment:10 Changed 9 years ago by mark1

Hyperspectral delivery created and ready for checking at:

~airborne/workspace/GB07_12-2010_082a_Debden/delivery/20100621/GB07_12/

comment:11 Changed 9 years ago by mark1

No Raw data for Hawk line "pre 5" but was reflown later in the flight as "pre 5R".

comment:12 Changed 9 years ago by knpa

Delivery checked.

All looks good apart from:

  • No bin dir listed in Readme contents.
  • Logsheet out of order in contents.
  • azgcorr command, new lines are not escaped.
  • azgcorr command, dem path pointing to "/users/rsg/arsf/dems/"?
  • Readme needs to be unix2dos'd
  • only 62 bands on eagle...correct?
  • Some dropped scan lines in hawk, should be mentioned in Readme
  • fastQC says hawk05 file size doesn't equal lines x samples x bands x 2 (and it's right). Mike says to process that line again?

comment:13 Changed 9 years ago by mark1

Above corrections have been applied.

comment:14 Changed 9 years ago by mark1

Eagle Line SCT Hawk Line SCT
-1 0.03 -1 0.0
-2 0.04 -2 0.02
-3 0.04 -3 0.04
-4 0.05 -4 -0.01
-5 0.71 -5 0.04
-6 0.05 -6 0.02
-7 0.02 -7 0.02
-8 0.05 -8 0.0
-9 0.04 -9 0.01
-10 0.03 N/A N/A

comment:15 Changed 9 years ago by mark1

LIDAR delivery directory has been created, and ready for checking, at:

~airborne/workspace/GB07_12-2010_082a_Debden/delivery/20100714/GB07_12

comment:16 Changed 9 years ago by chrfi

delivery checked, no problems found

comment:17 Changed 9 years ago by mark1

  • Description modified (diff)

Lidar and Photography dispatched to PI (23/07/2010)

comment:18 Changed 9 years ago by mark1

  • Description modified (diff)

Eagle data dispatched to PI on 07/09/2010

comment:19 Changed 9 years ago by mark1

rsyncing data, 8th September 2010, Hawk still needs to be processed so has not been rsynced.

comment:20 Changed 9 years ago by emca

Delivery check the Hawk data.
Needs bad pixel file added and flight letter added to file names (082 to be changed to 082a). Otherwise good to deliver.

comment:21 Changed 9 years ago by emca

Hawk data sent 08/12/2010

comment:22 Changed 9 years ago by emca

  • Description modified (diff)

comment:23 Changed 9 years ago by emca

Hawk data has been checked - ready for re-delivery

comment:24 Changed 9 years ago by emca

  • Description modified (diff)

Hawk data re-delivered 20/12/2010

comment:25 Changed 9 years ago by mark1

  • Component changed from Processing: general to Archiving

rsynced remainder of data (hawk delivery) 01/02/2011.
Ready for archiving.

comment:26 Changed 8 years ago by emca

Hyperspectral data reprocessed using apl software

No hawk data for line Pre 5, which is the third line.
Used numbers 1-10 for Eagle and 1-2, 4-10 for Hawk so the line number reflect the same flightline
Hawk raw file 3 is named line 4 in delivery, 4 is 5 etc.

No sct value for Eagle line 5, new sct value 0.71

Delivery sent to Martin Wooster 23/04/2012

comment:27 Changed 7 years ago by knpa

  • Owner changed from mark1 to knpa
  • Status changed from accepted to assigned

Beginning archiving.

comment:28 Changed 7 years ago by knpa

Made available to NEODC.

comment:29 Changed 6 years ago by knpa

Backed up by NEODC.

Moved to archive disk.

comment:30 Changed 6 years ago by knpa

  • Description modified (diff)

comment:31 Changed 3 years ago by dac

Archiving

Data available from NEODC (http://browse.ceda.ac.uk/browse/neodc/arsf/2010/GB07_12/GB07_12-2010_082a_Debden).

Marking as closed.

comment:32 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed
Note: See TracTickets for help on using tickets.